Knowledge Check 01 Helio Company has two products: A and B. The annual production and sales of Product A is 1,850 units and oActivities and Activity Measures Hand-processing (direct labor-hours) Setting up equipment (number of setups ) AdministeringKnowledge Check 01 Activities Assembly Processing Orders inspection Activity Rates $14.35 per machine-hour $ 47.85 per orderKnowledge Check 01 Harrington Company has two products: A and B. The annual production and sales of Product A is 1,750 units
Answer #1

1) Estimated labor hour = (1850*.30+1250*.60) = 1305 Hour

Predetermine overhead rate = Estimated overhead/estimated hour = 100485/1305 = 77 per hour

2) Hand processing activity rate = 320000/100000 = 3.20 per labor hour

So answer is d) $3.20 per labor hour

3) Calculate unit product cost

Assembly (690*14.35) 9901.50
Processing orders (40*47.85) 1914
Inspection (70.30*10) 703
Total overhead 12518.50
Overhead cost per unit (12518.50/430) 29.11
Direct material 35.82
Direct labor 29.56
Unit product cost 94.49

4) The amount of overhead cost that will be allocated to each unit of product B = 66*.70 = 46.20